# Efficient Inference in First Passage Time Models

### v1

@@ 6 Discussion
The benefits over existing approaches are especially notable for the likelihood computation of GDDMs with trial-specific dynamical covariates such as the attentional DDM (aDDM). Despite strong scientific interest, applications of aDDMs have been constrained by computational challenges. Current approaches to inference in aDDMs are based on replacing the aDDM with the gaze-weighted linear accumulator model (GLAM) [38, 40]. The GLAM replaces the time-varying drift rate $\mu(t)$ in a single trial of the aDDM with the constant drift rate $\tau^{-1}\int_{0}^{\tau}\mu(t)\,\mathrm{d}t$ that is simply the temporal average of the aDDM drift rate on that trial. Although the GLAM enables fast parameter inference using analytic formulas or simulation-based inference (SBI), it is not a well-specified statistical model, since the drift rate on a trial depends on the decision time on that trial. In parallel work [37], we show that maximum likelihood estimators based on the GLAM likelihood are not consistent for the corresponding aDDM parameters. This further illustrates the importance of developing practical computational tools for the complex GDDMs of interest in cognitive psychology and other fields.
@@
Our approach for multi-stage approximation should be generalizable to broader classes of SDEs than the ones considered here if they admit analytic formulas for FPTDs and NPDs in simple cases and if they have the appropriate Markov structure. If analytic formulas cannot be found, the multistage approximation might still be useful if accurate approximations of single-stage FPTDs and NPDs can be constructed by other methods, such as SBI.
There are several important mathematical questions that we leave for future work relating to regularity conditions for the existence of FPTDs, regularity conditions for the convergence of approximate multistage FPTDs under increasingly fine partitioning, and accuracy of numerical approximations for infinite series truncation, piecewise linear approximation of boundaries, and numerical quadrature. Despite these unresolved issues we believe our algorithms will find immediate use in the psychological and cognitive sciences and will improve the state-of-the-art for statistical inference of GDDMs.
Acknowledgment
This work was funded by National Institute of Mental Health grants P50 MH119467-01 and P50 MH106435-06A1, Office of Naval Research MURI Award N00014-23-1-2792, and was additionally supported by the Brainstorm Program at the Robert J. and Nancy D. Carney Institute for Brain Science.
Appendix A Proof of Almost Sure Convergence of First Passage Times under Refined Boundary Approximations
In this section, we establish the following result, which serves as the mathematical foundation of approximating the first passage times on curved boundaries by first passage times on their piece-wise linear approximations, as described in Section 4.4. Specifically, we consider the simplified setting where the diffusion process is a Brownian motion with a single upper boundary.
Theorem A.1.
Let $W(t)$ be a one-dimensional Brownian motion on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, and let $u(t)$ be a function on $[0,\infty)$ that is Hölder continuous with exponent $1/2$, satisfying $u(0)>0$. Suppose that a sequence of continuous functions $\{u_{n}\}$ converges uniformly to $u$ on compact subsets of $[0,\infty)$. Define the first passage times $\tau,\tau_{n}$ as
@@ Appendix A Proof of Almost Sure Convergence of First Passage Times under Refined Boundary Approximations
We prove the desired result in three steps.
  • $\liminf_{n\rightarrow\infty}\tau_{n}\geq\tau$.
    Fix $\omega\in\Omega$. For any $\epsilon>0$, the function $t\mapsto u(t)-W(t)$ is continuous and attain its minimum $\delta=\delta(\epsilon)>0$ on $[0,\tau-\epsilon]$, i.e., $W(t)<u(t)-\delta$ when $t\leq\tau-\epsilon$. Since $u_{n}\rightarrow u$ uniformly on $[0,\tau-\epsilon]$, there exists $N=N(\delta)=N(\epsilon)$ such that for all $n\geq N$ we have $u_{n}(t)>u(t)-\delta>W(t)$ for $t<\tau-\epsilon$. Thus, we have that for all $t<\tau-\epsilon$ and $n\geq N$, $W(t)<u_{n}(t)$. This implies $\tau_{n}\geq\tau-\epsilon$ for $n\geq N$. Taking $\epsilon\rightarrow 0$ gives $\liminf_{n\rightarrow\infty}\tau_{n}\geq\tau$.
  • $\limsup_{n\rightarrow\infty}\tau^{*}_{n}\leq\tau^{*}$.
    Fix $\omega\in\Omega$. By the definition of $\tau^{*}$, we know that for any $\epsilon>0$, there exists $t_{\epsilon}\in(\tau^{*},\tau^{*}+\epsilon)$ and $\delta>0$, such that $W(t_{\epsilon})>u(t_{\epsilon})+\delta$. Since $u_{n}(t_{\epsilon})\rightarrow u(t_{\epsilon})$, there exists $N=N(\delta)=N(\epsilon)$ such that for all $n\geq N$ we have $u_{n}(t_{\epsilon})<u(t_{\epsilon})+\delta<W(t_{\epsilon})$. Thus, there exists a $t_{\epsilon}\in(\tau^{*},\tau^{*}+\epsilon)$ such that $u_{n}(t_{\epsilon})<W(t_{\epsilon})$ for $n\geq N$. This implies $\tau^{*}_{n}\leq\tau^{*}+\epsilon$ for $n\geq N$. Taking $\epsilon\rightarrow 0$ gives $\limsup_{n\rightarrow\infty}\tau^{*}_{n}\leq\tau^{*}$. Hence $\limsup_{n\rightarrow\infty}\tau^{*}_{n}\leq\tau^{*}$.
  • $\tau=\tau^{*}$ a.s.
    The inequality $\tau\leq\tau^{*}$ is immediate, so it suffices to show that $\mathbb{P}\left(\tau<\tau^{*}\right)=0$. Without loss of generality, we assume that $\tau$ is almost surely finite, since for any $\omega$ where $\tau(\omega)=\infty$, we trivially have $\tau(\omega)=\tau^{*}(\omega)=\infty$.
@@
    $\limsup_{s\rightarrow 0^{+}}\frac{B(s)}{\sqrt{2s\log\log(1/s)}}\leq\limsup_{s\rightarrow 0^{+}}\frac{C}{\sqrt{2\log\log(1/s)}}=0$ (22)
@@
[…]
    so equation (22) holds with probability 00. So $\mathbb{P}(\tau<\tau^{*})=0$.
Combining the above results, we obtain
[…]
@@
Adopting the notations in Section 4.4 and the footnote in Section 2, the piecewise linear interpolations of the continuous boundaries $u$ and $\ell$, denoted by $\bar{u}$ and $\bar{\ell}$ respectively, converge to $u$ and $\ell$ uniformly on compact subsets of $[0,\infty)$ as the interpolation grid is refined (see P35 of [12]). Consequently, we can invoke Theorem A.1 to conclude that the approximate first passage time $\bar{\tau}_{u}$ and $\bar{\tau}_{\ell}$ converge almost surely to $\tau_{u}$ and $\tau_{\ell}$ on their respective boundaries. Since $\bar{\tau}=\bar{\tau}_{u}\wedge\bar{\tau}_{\ell},\tau=\tau_{u}\wedge\tau_{\ell}$, it follows that $\bar{\tau}\rightarrow\tau$ a.s..
Appendix B Transformation of a Diffusion Process to a Brownian Motion
@@
Theorem B.1 ([53]).
Let $X(t)$ be a diffusion process defined on interval $[0,T_{\text{end}}]$ satisfying the $\operatorname{SDE}$ (2). If there exists a pair of functions $c_{1}(t)$ and $c_{2}(t)$ such that
$\mu(x,t)=\frac{\frac{\partial}{\partial x}\left(\sigma^{2}(x,t)\right)}{4}+\frac{\sigma(x,t)}{2}\Big{(}c_{1}(t)+\int_{0}^{x}\frac{c_{2}(t)\sigma^{2}(y,t)+\frac{\partial}{\partial t}\left(\sigma^{2}(y,t)\right)}{\sigma^{3}(y,t)}\,\mathrm{d}y\Big{)}$
@@
$w$ $=\psi(x,t)=e^{-\frac{1}{2}\int_{0}^{t}c_{2}(s)\,\mathrm{d}s}\int_{0}^{x}\frac{1}{\sigma(y,t)}\,\mathrm{d}y-\frac{1}{2}\int_{0}^{t}c_{1}(r)e^{-\frac{1}{2}\int_{0}^{r}c_{2}(z)\,\mathrm{d}z}\,\mathrm{d}r$
We take the Ornstein-Uhlenbeck process (18) as an example. Here $\mu(x,t)=\theta(\lambda-x)$ and $\sigma(x,t)=\sigma$, and we can take $c_{1}(t)=\frac{2\theta\lambda}{\sigma},c_{2}(t)=-2\theta$. Correspondingly, the transformation is given by $s=\gamma(t)=\frac{e^{2\theta t}-1}{2\theta},w=\psi(x,t)=\frac{1}{\sigma}(e^{\theta t}x-\lambda e^{\theta t}+\lambda)$. This is the same as the transformation given in Section 4.5.
Appendix C Detailed Implementation of the Quadrature
The $m$-point Gauss-Legendre quadrature is given by
where $\mathbf{w}=(w_{1},\cdots,w_{m})^{\top}$ are the quadrature weights, $\boldsymbol{\xi}=(\xi_{1},\cdots,\xi_{m})^{\top}$ are the quadrature points on the reference interval $[-1,1]$. The formula (23) is exact for polynomials of degree up to $12m-$ 1. $\mathbf{w}$ and $\boldsymbol{\xi}$ can be computed by formulas involving Legendre polynomials but are usually tabulated to great accuracy.
To apply formula (23) to our algorithm, at each stage, we only need to track the density at precomputed quadrature points on each vertical boundary. We then proceed to the next stage by computing the transitional density to the quadrature points on the next vertical boundary (see Figure 8). Suppose we use quadrature of order $m_{k}$ for integrating in the $k$-th stage. The functions in equations (6) and (7) are observed to exhibit a more rapid growth of derivatives compared to those in equation (8), hence the integrand in the final stage is less regular than in the earlier stages. Therefore, we typically employ a higher-order quadrature rule for the integration in the final stage. Specifically, if $\tau$ is in the $K$-th stage we can take $m_{K}\geq m_{i}$ for $i=1,\cdots,K-1$ for more efficient computation.
Here, we present Algorithm 2, which incorporates numerical integration via (23) into Algorithm 1. Note that different initial treatments are needed for discrete and continuous $X(0)$, as if $X(0)$ is discrete, we only need to directly use formulas (9), (10) and (11), and there is no need for integration in the first stage. In cases where $X(0)$ is a mixture of discrete and continuous components, we handle each part accordingly. However, we omit this case in Algorithm 2 for brevity, as its treatment follows naturally from the discrete and continuous cases. In Algorithm 2, we denote the abstract process of acquiring $\mathbf{w},\boldsymbol{\xi}$ of order $m$ as $\mathbf{w},\boldsymbol{\xi}\leftarrow\text{gausslegendre}(m)$, and the symbol $\odot$ represents the Hadamard (element-wise) product for vectors.
@@ Appendix C Detailed Implementation of the Quadrature
if $q_{0}\text{ is a pdf}$ then /* $X(0)$ is continuous */

### v2

@@ 6 Discussion
The benefits over existing approaches are especially notable for the likelihood computation of GDDMs with trial-specific covariates such as the aDDM. Despite strong scientific interest, applications of aDDMs have been constrained by computational challenges. Current approaches to inference in aDDMs rely on unjustified computational simplifications, such as the gaze-weighted linear accumulator model (GLAM) [35, 37]. The GLAM replaces the time-varying drift rate $\mu(t)$ in a single trial of the aDDM with the constant drift rate $\tau^{-1}\int_{0}^{\tau}\mu(t)\,\mathrm{d}t$ that is simply the temporal average of the aDDM drift rate on that trial. Although the GLAM enables fast parameter inference using analytic formulas, it is not a well-specified statistical model, since the drift rate on a trial depends on the decision time on that trial. In parallel work [34], we show that maximum likelihood estimators based on the GLAM likelihood are not consistent for the corresponding aDDM parameters. This further illustrates the importance of developing practical computational tools for the complex GDDMs of interest in cognitive neuroscience and other fields.
@@
Our approach for multi-stage approximation should be generalizable to broader classes of SDEs than the ones considered here if they admit analytic formulas for FPTDs and NPDs in simple cases and if they have the appropriate Markov structure. If analytic formulas cannot be found, the multi-stage approximation might still be useful if accurate approximations of single-stage FPTDs and NPDs can be constructed by other methods, such as SBI.
There are several important mathematical questions that we leave for future work relating to regularity conditions for the existence of FPTDs, regularity conditions for the convergence of approximate multi-stage FPTDs under increasingly fine partitioning, and accuracy of numerical approximations for infinite series truncation, piecewise linear approximation of boundaries, and numerical quadrature. Despite these unresolved issues we believe our algorithms will find immediate use in the psychological and cognitive sciences and will improve the state-of-the-art for statistical inference of GDDMs.
Declarations
[…]
Acknowledgements This work was supported by the National Institute of Mental Health (grants P50 MH119467-01 and P50 MH106435-06A1), the Office of Naval Research (MURI Award N00014-23-1-2792), and the Brainstorm Program at the Robert J. and Nancy D. Carney Institute for Brain Science. S.L. was additionally supported by the JBB Endowed Graduate Fellowship in Brain Science.
[…]
Contributions M.J.F. and A.F. raised the research question and provided expertise on the computational cognitive science applications. M.T.H. and S.L. developed the methodological and theoretical framework. S.L. carried out the research, validated the results, and prepared the associated GitHub repository with support from all authors. S.L. wrote the original draft with feedback and revisions from all authors. M.J.F. secured funding for the project. M.T.H supervised the project.
[…]
Data Availability
The code supporting this study will be released as a package and made publicly available at https://github.com/RiverFlowsInYou98/efficient-fpt upon publication.
[…]
Competing interests
The authors declare no competing interests.
Appendix A
A.1 Proof of Almost Sure Convergence of First Passage Times under Refined Boundary Approximations
In this section, we establish a continuity result, which serves as the mathematical foundation of approximating the first passage times on curved boundaries by first passage times on their piecewise linear approximations, as described in Section 4.4. We first consider the simplified setting of a Brownian motion hitting a single upper boundary.
Theorem 1.
Let $W(t)$ be a one-dimensional Brownian motion on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, and let $u(t)$ be a function on $0,\infty)$ that is Hölder continuous with exponent $1/2$, satisfying $u(0)>0$. Suppose that a sequence of continuous functions $(u_{n})$ converges uniformly to $u$ on compact subsets of $0,\infty)$. Define the first passage times $\tau,\tau_{n}$ as
@@ A.1 Proof of Almost Sure Convergence of First Passage Times under Refined Boundary Approximations
We prove the desired result in three steps. We first prove that $\liminf_{n\rightarrow\infty}\tau_{n}\geq\tau$. Fix $\omega\in\Omega$. For any $\epsilon>0$, the function $t\mapsto u(t)-W(t)$ is continuous and attain its minimum $\delta=\delta(\epsilon)>0$ on $0,\tau-\epsilon]$, i.e., $W(t)<u(t)-\delta$ when $t\leq\tau-\epsilon$. Since $u_{n}\rightarrow u$ uniformly on $0,\tau-\epsilon]$, there exists $N=N(\delta)=N(\epsilon)$ such that for all $n\geq N$ we have $u_{n}(t)>u(t)-\delta>W(t)$ for $t<\tau-\epsilon$. Thus, we have that for all $t<\tau-\epsilon$ and $n\geq N$, $W(t)<u_{n}(t)$. This implies $\tau_{n}\geq\tau-\epsilon$ for $n\geq N$. Taking $\epsilon\rightarrow 0$ gives $\liminf_{n\rightarrow\infty}\tau_{n}\geq\tau$.
Next, we prove that $\limsup_{n\rightarrow\infty}\tau^{*}_{n}\leq\tau^{*}$. Fix $\omega\in\Omega$. By the definition of $\tau^{*}$, we know that for any $\epsilon>0$, there exists $t_{\epsilon}\in(\tau^{*},\tau^{*}+\epsilon)$ and $\delta>0$, such that $W(t_{\epsilon})>u(t_{\epsilon})+\delta$. Since $u_{n}(t_{\epsilon})\rightarrow u(t_{\epsilon})$, there exists $N=N(\delta)=N(\epsilon)$ such that for all $n\geq N$ we have $u_{n}(t_{\epsilon})<u(t_{\epsilon})+\delta<W(t_{\epsilon})$. Thus, there exists a $t_{\epsilon}\in(\tau^{*},\tau^{*}+\epsilon)$ such that $u_{n}(t_{\epsilon})<W(t_{\epsilon})$ for $n\geq N$. This implies $\tau^{*}_{n}\leq\tau^{*}+\epsilon$ for $n\geq N$. Taking $\epsilon\rightarrow 0$ gives $\limsup_{n\rightarrow\infty}\tau^{*}_{n}\leq\tau^{*}$. Hence $\limsup_{n\rightarrow\infty}\tau^{*}_{n}\leq\tau^{*}$.
Finally, we prove that $\tau=\tau^{*}$ a.s.. The inequality $\tau\leq\tau^{*}$ is immediate, so it suffices to show that $\mathbb{P}\left(\tau<\tau^{*}\right)=0$. Without loss of generality, we assume that $\tau$ is almost surely finite, since for any $\omega$ where $\tau(\omega)=\infty$, we trivially have $\tau(\omega)=\tau^{*}(\omega)=\infty$.
@@
$\limsup_{s\rightarrow 0^{+}}\frac{B(s)}{\sqrt{2s\log\log(1/s)}}$ (22)
$\leq$ $\limsup_{s\rightarrow 0^{+}}\frac{C}{\sqrt{2\log\log(1/s)}}=0$
@@
$\limsup_{s\rightarrow 0^{+}}\frac{B(s)}{\sqrt{2s\log\log(1/s)}}=1\quad\text{a.s.}$
so equation (22) holds with probability 0. So $\mathbb{P}(\tau<\tau^{*})=0$.
Combining the above three steps, we obtain
$\tau\leq\liminf_{n\rightarrow\infty}\tau_{n}\leq\limsup_{n\rightarrow\infty}\tau_{n}\leq\limsup_{n\rightarrow\infty}\tau_{n}^{*}\leq\tau^{*}$
@@
Returning to the two-boundary setting, and adopting the notation of Section 4.4, the piecewise linear interpolations of the Hölder-$1/2$ continuous boundaries $u$ and $\ell$, denoted by $\overline{u}$ and $\overline{\ell}$ respectively, converge to $u$ and $\ell$ uniformly on compact subsets of $0,\infty)$ as the interpolation grid is refined (see p35 of [9]). Consequently, let $\bar{\tau}_{u}$ and $\bar{\tau}_{\ell}$ denote the approximate one-sided first passage times, and $\tau_{u}$ and $\tau_{\ell}$ their exact counterparts, then we can invoke Theorem 1 to conclude that the $\overline{\tau}_{u}$ and $\overline{\tau}_{\ell}$ converge almost surely to $\tau_{u}$ and $\tau_{\ell}$ on their respective boundaries. Since $\overline{\tau}=\overline{\tau}_{u}\wedge\overline{\tau}_{\ell},\tau=\tau_{u}\wedge\tau_{\ell}$, it follows that $\overline{\tau}\rightarrow\tau$ a.s.
A.2 Transformation of a Diffusion Process to a Brownian Motion
@@
Theorem 2 ([47]).
Let $X(t)$ be a diffusion process defined on interval $0,T_{\text{end}}]$ satisfying the $\operatorname{SDE}$ (2). If there exists a pair of functions $c_{1}(t)$ and $c_{2}(t)$ such that
$\mu(x,t)=\frac{\frac{\partial}{\partial x}\left(\sigma^{2}(x,t)\right)}{4}+$ (23)
$\quad\frac{\sigma(x,t)}{2}\Big(c_{1}(t)+\int_{0}^{x}\frac{c_{2}(t)\sigma^{2}(y,t)+\frac{\partial}{\partial t}\left(\sigma^{2}(y,t)\right)}{\sigma^{3}(y,t)}\,\mathrm{d}y\Big)$
@@
$w=\psi(x,t)=$ $e^{-\frac{1}{2}\int_{0}^{t}c_{2}(s)\,\mathrm{d}s}\int_{0}^{x}\frac{1}{\sigma(y,t)}\,\mathrm{d}y$
$-\frac{1}{2}\int_{0}^{t}c_{1}(r)e^{-\frac{1}{2}\int_{0}^{r}c_{2}(z)\,\mathrm{d}z}\,\mathrm{d}r$
We take the Ornstein-Uhlenbeck process (18) as an example. Here $\mu(x,t)=\theta(\lambda-x)$ and $\sigma(x,t)=\sigma$, and the Cherkasov condition (23) reads
$\theta(\lambda-x)=\frac{\sigma}{2}c_{1}(t)+\frac{x}{2}c_{2}(t)$
We can take $c_{1}(t)=\frac{2\theta\lambda}{\sigma},c_{2}(t)=-2\theta$. Correspondingly, the transformation is given by $s=\gamma(t)=\frac{e^{2\theta t}-1}{2\theta},w=\psi(x,t)=\frac{1}{\sigma}(e^{\theta t}x-\lambda e^{\theta t}+\lambda)$. This is the same as the transformation given in Section 4.5. Verifying the Cherkasov condition for general diffusion processes requires analyzing Eq. (23) on a case-by-case basis.
A.3 Fast Implementation
In this section, we discuss some practical aspects of our algorithm that contribute to its fast execution, including important details about numerical integration and function representation that are glossed over in the description of Algorithm 1. We present Algorithm 2, the practical version of Algorithm 1 to conclude this section.
To compute the likelihood given a dataset, we simply iterate Algorithm 1 over all data. Additional dataset-level speedups can be achieved using the following strategies:
  • When multiple observations share the same GDDM configuration, a single pass of Algorithm 1 suffices to evaluate their likelihoods simultaneously, eliminating redundant computations. This idea underlies the speedups achieved by many previous methods (e.g., KFE-based approaches). However, such methods slow down considerably when GDDMs vary across trials, whereas our approach adapts seamlessly to both homogeneous and heterogeneous settings while maintaining high computational efficiency.
  • The computation of each configuration in a dataset is independent of each other, rendering the problem “embarrassingly parallelizable” across configurations — we can distribute the dataset over multiple threads, and since no communication is needed between the sub-tasks, the computational time should ideally be inversely proportional to the number of threads, up to the point where the number of threads equals the number of configurations. Consequently, models such as the aDDM, where each trial has a distinct configuration, can benefit substantially from this form of parallelization. We repeat the timing experiment for Algorithm 1 described in Section 5.4.1 and report the results in Table 3, which demonstrate that our method can easily be further accelerated through parallelization.
Table 3: Evaluation times for the aDDM likelihood (21) with 50,00050{,}000 trial data, using Algorithm 1 under different thread counts.
# Threads 1 2 4 8 16 32
Time(s)($\mathrm${s}) 8.393 4.220 2.130 1.087 0.565 0.304
A.3.1 Detailed Implementation of the Quadrature
@@
1exif $q_{0}\text{ is a pdf}$ then /* $X(0)$ is continuous */
@@
The $m$-point Gauss-Legendre quadrature is given by
where $\mathbf{w}=(w_{1},\cdots,w_{m})^{\top}$ are the quadrature weights, $\boldsymbol{\xi}=(\xi_{1},\cdots,\xi_{m})^{\top}$ are the quadrature points on the reference interval [1,1][-1,1]. The formula (24) is exact for polynomials of degree up to 2m12m-1. $\mathbf{w}$ and $\boldsymbol{\xi}$ can be computed by formulas involving Legendre polynomials but are usually tabulated to great accuracy.
To apply formula (24) to our algorithm, at each stage $k$, we only need to track the density at precomputed quadrature points on each vertical boundary. We then proceed to the next stage by computing the transitional densities $q_{k}$ from the current quadrature points to the quadrature points on the next vertical boundary (see Figure 9). Regarding the order of quadrature, the functions in equations (6) and (7) are observed to exhibit a more rapid growth of derivatives compared to those in equation (8), hence the integrand in the final stage is less regular than in the earlier stages. Therefore, we typically employ a quadrature rule of equal or higher order in the final stage. In other words, if we let $m_{k}$ denote the quadrature order for integration in the $k$-th stage, then when $\tau$ lies in the $K$-th stage, we can choose $m_{K}\geq m_{i}$ for $i=1,\ldots,K-1$ to achieve more efficient computation.
Here, we present Algorithm 2, which incorporates numerical integration via (24) into Algorithm 1. Note that different initial treatments are needed for discrete and continuous $X(0)$, as if $X(0)$ is discrete, we only need to directly use formulas (4.1), (4.1) and (4.1), and there is no need for integration in the first stage. In cases where $X(0)$ is a mixture of discrete and continuous components, we handle each part accordingly. However, we omit this case in Algorithm 2 for brevity, as its treatment follows naturally from the discrete and continuous cases.
In Algorithm 2, we denote the abstract process of acquiring $\mathbf{w},\boldsymbol{\xi}$ of order $m$ as $\mathbf{w},\boldsymbol{\xi}\leftarrow\text{gausslegendre}(m)$, and the symbol $\odot$ represents the Hadamard (element-wise) product for vectors. Assuming constant time complexity for evaluating the single-stage FPTDs $f_{u}^{\text{single}},f_{\ell}^{\text{single}}$ and NPD $q^{\text{single}}$, the overall time complexity of Algorithm 2 is $\Theta(\sum_{k=1}^{K}m_{k-1}m_{k})$.